# A Survey on the Application of Large Language Models in Scenario-Based Testing of Automated Driving Systems

### v1

@@ Abstract
The safety and reliability of Automated Driving Systems (ADSs) must be validated prior to large-scale deployment. Among existing validation approaches, scenario-based testing has been regarded as a promising method to improve testing efficiency and reduce associated costs. Recently, the emergence of Large Language Models (LLMs) has introduced new opportunities to reinforce this approach. While an increasing number of studies have explored the use of LLMs in the field of automated driving, a dedicated review focusing on their application within scenario-based testing remains absent. This survey addresses this gap by systematically categorizing the roles played by LLMs across various phased of scenario-based testing, drawing from both academic research and industrial practice. In addition, key characteristics of LLMs and corresponding usage strategies are comprehensively summarized. The paper concludes by outlining five open challenges and potential research directions. To support ongoing research efforts, a continuously updated repository of recent advancements and relevant open-source tools is made available at: https://github.com/ftgTUGraz/LLM4ADSTest.
Index Terms:
Generative AI, simulation test, safety assessment, automated vehicle, literature review.I Introduction
@@
Before the deployment of Automated Driving Systems (ADSs), their safety and reliability must be rigorously validated, a process traditionally requiring billions of miles on-road driving by Automated Vehicles (AVs) [1]. However, conventional mileage-based on-road testing has been deemed impractical due to its high costs and time-intensive nature. To address these limitations, the PEGASUS project [2] introduced the scenario-based testing approach, which enhances validation efficiency by subjecting ADSs to simulated and virtual driving environments. The effectiveness of this method has been substantiated, with simulation-based testing recongnized as a major contributor to the performance improvements of AVs by Waymo [3]. Consequently, the exploration of scenario-based testing has emerged as a crucial area of research.
@@
Meanwhile, rapid advancements have been observed in the field of Artificial Intelligence (AI). Among the most prominent developments, Large Language Models (LLMs) have emerged as state-of-the-art AI systems capable of understanding and generating human language, and of performing a wide range of tasks without task-specific training [4, 5]. In recent years, researches at the intersection of LLMs and Automated Driving (AD) have gained increasing attention. As shown in Figure 1, subplot 1a illustrates the increasing volume of research publications related to LLMs and AD topics individually, while subplot 1b highlights a noticeable rise in studies specifically addressing the application of LLMs in the AD domain. This trend underscores a rapidly growing research direction focused on applying LLMs in development of ADSs.
@@
Existing surveys have reviewed various aspects of LLMs, including their development process [6, 7, 8], reasoning capabilities [9] and prompting strategies [10, 11], as well as privacy-related concerns [12, 13, 14]. In parallel, several surveys have addressed scenario-based testing methodologies [15, 16, 17, 18, 19, 20, 21, 22, 23], and others have examined the application of LLMs in AD and related domains [24, 25, 26, 27, 28]. However, a systematic review specifically focusing on how LLMs are utilized within scenario-based testing of ADSs remains absent.
@@
The remainder of this work is organized as follows: Section II reviews recent survey studies categorized by relevant topics. Section III introduces key terminologies used throughout the paper. Section IV presents a detailed analysis of specific tasks performed by LLMs. Section V summarizes the types of LLMs and their corresponding adoption strategies using tabular representations. Section VI briefly outlines the application of LLMs in industrial contexts. Section VII identifies five major challenges and discusses potential research directions. Finally, Section VIII presents the concluding remarks of this survey.
II Related Survey
This section provides an overview of recent relevant survey studies. It is organized into four categories: general surveys on LLMs, surveys focusing on scenario-based testing, surveys covering the use of LLMs in AD, and surveys addressing LLM applications in other domains.
II-A Survey of LLM
Recent surveys have examined the development and evolution of LLMs and their multimodal extensions. Naveed et al. [6] and Hadi et al. [7] offer comprehensive overviews of LLM architectures, training strategies, practical applications, and associated challenges. Zhang et al [8] systematically analyze 126 multimodal LLMs (MLLMs), with detailed comparison of model design and benchmark performances.
In addition to model development, several surveys have explored key functional aspects of LLMs. Su et al. [9] review efficient reasoning strategies, categorizing them into model-based, output-based, and prompt-based approaches. The domain of prompt engineering has been addressed by Sahoo et al. [10] and Chen et al. [11], who examine various methods, datasets, applications, and associated security implications.
Privacy preservation remains a critical limitation to broader deployment of LLMs. To address this concern, multiple surveys have investigated privacy-related challenges. Hu et al. [12] discuss the application of differential privacy in natural language models. Neel and Chang [13] provide a broader overview of privacy issues in LLMs, while Edemacu and Wu [14] focus specifically on privacy concerns related to in in-context learning and prompting mechanisms.
II-B Survey of Scenario-Based Testing
Scenario-based testing has been widely recognized as a critical methodology for assessing the safety and reliability of ADSs. A number of studies have reviewed this approach from various perspectives. Nalic et al. [15] and Zhang et al. [16] conduct extensive reviews encompassing over 80 publications and proposed taxonomies to classify scenario generation techniques. Zhong et al. [17] focus on scenario-based testing utilizing high-fidelity simulation platforms, whereas Ding et al. [18] categorize generation methods into data-driven, adversarial, and knowledge-based categories.
Beyond scenario generation, some studies have extended their scope to include safety assessment frameworks. Riedmaier et al. [19] advocate for the integration of formal verification methods within scenario-based testing. Similarly, Wishart et al. [20] review current verification and validation practices to support the development of SAE standards. Further, Alghodhaifi et al. [21] and Tang et al [22] investigate evaluation efficiency, emphasizing the role of accelerated and AI-based techniques as valuable complements to scenario-based approaches. Finally, Mihalj et al. [23] explore the interaction between ADSs and physical infrastructure, offering perspectives on how environmental factors can influence the design of test scenarios.
II-C Survey of LLM for Automated Driving
Recent surveys investigate the integration of LLMs and their multimodal extensions into ADSs, highlighting their potential to enhance perception, reasoning, and decision-making capabilities. Within a particular focus on Vision-Language Models (VLMs), Yang et al. [29] and Zhou et al. [30] review their applications across key functions such as perception, planning, control, and data generation, while underscoring their open-world understanding capabilities and the challenges involved. In a complementary study, Cui et al. [31] systematically review the development, tools, datasets, and challenges of applying MLLMs in AD and map systems, with the aim of informing future research and practical implementation.
Broader discussions are provided by Fourati et al. [32] and Li et al. [33], who examine the deployment of LLMs, VLMs and MLLMs within both modular and end-to-end system architecture, focusing on their structural designs, deployment strategies, and prospective research directions. At the system-level, Gan et al. [34] offer a comprehensive review of LLM applications in intelligent transportation systems, with an emphasis on scalable and efficient implementation.
II-D Survey of LLM for Miscellaneous Domains
Beyond the domain of AD, LLMs are extensively reviewed across a range of application areas. In the context of software testing, Braberman et al. [24] propose a taxonomy for LLM-based verification tasks, while Zhou et al. [25] outline frameworks and challenges associated with intelligent system testing. Yu et al. [26] survey alignment algorithms for MLLMs, and Luo et al. [27] review visual foundation models for road scene understanding. In the context of urban innovation, Xu et al. [28] examine the integration of generative AI models with urban digital twins, emphasizing their potential to support smart city management.
III Terminology
III-A Scenario
III-B Scenario Abstraction Level

### v2

@@ Abstract
The safety and reliability of Automated Driving Systems (ADSs) must be validated prior to large-scale deployment. Among existing validation approaches, scenario-based testing has been regarded as a promising method to improve testing efficiency and reduce associated costs. Recently, the emergence of Large Language Models (LLMs) has introduced new opportunities to reinforce this approach. While an increasing number of studies have explored the use of LLMs in the field of automated driving, a dedicated review focusing on their application within scenario-based testing remains absent. This survey addresses this gap by systematically categorizing the roles played by LLMs across various phases of scenario-based testing, drawing from both academic research and industrial practice. In addition, key characteristics of LLMs and corresponding usage strategies are comprehensively summarized. The paper concludes by outlining five open challenges and potential research directions. To support ongoing research efforts, a continuously updated repository of recent advancements and relevant open-source tools is made available at: https://github.com/ftgTUGraz/LLM4ADSTest.
I Introduction
@@
Before the deployment of Automated Driving Systems (ADSs), their safety and reliability must be rigorously validated, a process traditionally requiring billions of miles of on-road driving by Automated Vehicles (AVs) [kalra2016driving]. However, conventional mileage-based on-road testing has been deemed impractical due to its high costs and time-intensive nature. To address these limitations, several initiatives, such as the PEGASUS project [winner2019pegasus], ENABLE-S3 [ENA2019], HEADSTART [HEADSTART_CORDIS], StreetWise [TNO_StreetWise], and UNECE WP29 [UNECE2023NATM] introduced the scenario-based testing approach, which enhances validation efficiency by subjecting ADSs to simulated and virtual driving environments. The effectiveness of this method has been substantiated, with simulation-based testing recognized as a major contributor to the performance improvements of AVs by Waymo [10.1145/3177753]. Since then, the exploration of scenario-based testing has emerged as a crucial area of research.
@@
Meanwhile, rapid advancements have been observed in the field of Artificial Intelligence (AI). Among the most prominent developments, Large Language Models (LLMs)111In this work, the term “LLM” is used as a convenient general reference, encompassing its multimodal extensions when discussing general application strategies. have emerged as state-of-the-art AI systems capable of understanding and generating human language, and of performing a wide range of tasks without task-specific training [radford2019language, brown2020language]. In recent years, research at the intersection of LLMs and Automated Driving (AD) have gained increasing attention. As shown in Figure 1, subplot 1a illustrates the increasing volume of research publications related to LLMs and AD topics individually, while subplot 1b highlights a noticeable rise in studies specifically addressing the application of LLMs in the AD and transportation domain. This trend underscores a rapidly growing research direction focused on applying LLMs in the development of ADSs.
@@
Existing surveys have reviewed various aspects of LLMs, including their development process [Naveed.2024, hadi2023survey, Zhang.2024], reasoning capabilities [sui2025stop] and prompting strategies [sahoo2024systematic, chen2023unleashing], as well as privacy-related concerns [hu2023differentially, neel2023privacy, edemacu2024privacy]. In parallel, several surveys have addressed scenario-based testing methodologies [Nalic.2020, zhang2022finding, Zhong.2021, Ding.2023, Riedmaier.2020, Wishart.2020, Alghodhaifi.2021, Tang.2023, mihalj2022road], and others have examined the application of LLMs in AD and related domains [Braberman.2024, Zhou.2024, yu2025aligning, Luo.2024, Xu.2024b]. However, a systematic review specifically focusing on how LLMs are utilized within scenario-based testing of ADSs remains absent.
@@
To facilitate community research, a continuously updated repository is maintained to provide a systematic bibliography, taxonomy tags, and evaluation tools to track rapid advancements in the field. The remainder of this work is organized as follows: Section II introduces fundamental terminologies used throughout the paper and reviews recent survey studies categorized by relevant topics. Section III presents a detailed analysis of specific tasks performed by LLMs. Section IV summarizes the types of LLMs and their corresponding adoption strategies using tabular representations. Section V briefly outlines the application of LLMs in industrial contexts. Section VI identifies five major challenges and discusses potential research directions. Finally, Section VII presents the concluding remarks of this survey.
II Fundamental Terminology and Related Survey
II-A Fundamental Terminology
II-A1 Acronym
To facilitate understanding of the technical terminology used throughout this survey, key acronyms are listed with their full terms in Table I.
| Acronym | Full Term |
| ADE | Average Displacement Error |
| ADS | Automated Driving System |
| AEB | Autonomous Emergency Braking |
| AI | Artificial Intelligence |
| API | Application Programming Interface |
| ASIL | Automotive Safety Integrity Level |
| AV | Automated Vehicle |
| BEV | Bird’s Eye View |
| CoT | Chain-of-Thought |
| CR | Collision Rate |
| DSL | Domain-Specific Language |
| HARA | Hazard Analysis and Risk Assessment |
| LLM | Large Language Model |
| LoRA | Low-Rank Adaptation |
| LVLM | Large Vision-Language Model |
| MLLM | Multimodal Large Language Model |
| NHTSA | National Highway Traffic Safety Administration |
| NL | Natural Language |
| NLP | Natural Language Processing |
| ODD | Operational Design Domain |
| OOD | Out-Of-Distribution |
| OS | Overall Score |
| PPO | Proximal Policy Optimization |
| RAG | Retrieval-Augmented Generation |
| RL | Reinforcement Learning |
| SQL | Structured Query Language |
| STPA | Systems Theoretic Process Analysis |
| TTC | Time-To-Collision |
| UNECE | United Nations Economic Commission for Europe |
| VLM | Vision-Language Model |
| WOD | Waymo Open Dataset |
| WOMD | Waymo Open Motion Dataset |
| XiL | X-in-the-Loop |
| XOSC | ASAM OpenSCENARIO |
| XODR | ASAM OpenDRIVE |
II-A2 Scenario
A scenario is a description of the temporal relationship between several scenes in a sequence of scenes, with goals and values within a specified situation, influenced by actions and events [ulbrich2015defining, ISO21448].
II-A3 Scenario Abstraction Level